Some nations require their students to pass an exam before earning their primary school degrees or diplomas. A certain nation gives students an exam whose scores are normally distributed with a mean of 41 points and a standard deviation of 9 points.

Suppose we select 2 of these testers at random, and define the random variable D as the difference between their scores.

Find the probability that their scores are within 10 points of each other.
You may round your answer to two decimal places.

P(∣D∣<10)≈
Answer #1

solution:- D=x,-X2 les: Elx,) - E(X) = 41-41 = 0 o = √V (x) + X(X) = √get g2 = ore Now, ze D-Up 12 P(IDI (10) = Pl-loc Dcio)